There is a class. The ratio is 5 boys to 4 girls. There are 65 boys. How many girls are there?
Answer by nerdybill(5404)

There is a class. The ratio is 5 boys to 4 girls. There are 65 boys. How many girls are there?
.
Let g = number of girls in the class
.
4:5 = g:65
or
4/5 = g/65
g = (4*65)/5
g = 260/5
g = 52 (girls in the class)